OpenStudy (noelgreco):

A class is given a list of 20 study problems from which 10 will be part of an upcoming exam. If a given student knows how to solve 15 of the problems, find the probability that the student will be able to answer: a. all ten questions on the exam b. exactly eight questions on the exam c. at least 9 questions on the exam

OpenStudy (anonymous):

I think this can be considered as a Bernoulli Trial. In that case, you could say that the student is in p=3/4 right and in p=1-p=1/4 false. This are of course only initial conditions, if I understand the problem correctly.

OpenStudy (noelgreco):

For a. I got (15C10)/20C10 b. (15C8)X(5C2)/(20C10) Both are correct. C has me stumped.